# When Machines Get It Wrong: Large Language Models Perpetuate Autism Myths More Than Humans Do

### v1

@@ Abstract
As Large Language Models become ubiquitous sources of health information, understanding their capacity to accurately represent stigmatized conditions is crucial for responsible deployment. This study examines whether leading AI systems perpetuate or challenge misconceptions about Autism Spectrum Disorder, a condition particularly vulnerable to harmful myths. We administered a 30-item instrument measuring autism knowledge to 178 Spanish participants and three state-of-the-art LLMs including GPT-4, Claude, and Gemini. Contrary to expectations that AI systems would leverage their vast training data to outperform humans, we found the opposite pattern: human participants endorsed significantly fewer myths than LLMs (36.2% vs. 44.8% error rate; z=2.59z=-2.59, $p=.0048$). In 18 of the 30 evaluated items, humans significantly outperformed AI systems. These findings reveal a critical blind spot in current AI systems and have important implications for human-AI interaction design, the epistemology of machine knowledge, and the need to center neurodivergent perspectives in AI development.
@@ 1 Introduction
This study provides the first systematic comparison of autism knowledge between leading LLMs and a human population sample. We designed a 30-item instrument capturing both myths and facts about autism, administered it to 178 Spanish participants, and posed identical items to GPT-4, Claude, and Gemini. Our investigation examined whether LLMs demonstrate greater or lesser accuracy than humans in distinguishing autism myths from facts, whether systematic differences exist in human-AI performance gaps across categories of autism knowledge, and what human characteristics predict better autism knowledge. The findings reveal a counterintuitive pattern: despite their vast training corpora, state-of-the-art LLMs endorse autism myths at significantly higher rates than human respondents.
@@ 3 Method
The autism knowledge instrument was developed through a systematic process designed to ensure content validity. Initial items were generated through literature review, extracting autism-related claims from peer-reviewed research on misconceptions [20, 5, 6], DSM-5 diagnostic criteria [10], and Spanish institutional documents [17, 19]. The final instrument contained 30 items comprising 16 myths and 14 facts. Respondents indicated agreement on a 5-point scale ranging from strongly disagree to strongly agree, and an attention check question was included to filter unreliable responses.
@@
Three leading LLMs were selected based on market prominence: GPT-4 from OpenAI, Claude from Anthropic, and Gemini from Google. API access was obtained through commercial subscriptions. Python scripts were developed using the official APIs to automate requests. Each of the 30 statements was sent to each model, and responses were recorded and transformed into proportions of agreement expressed as percentages. Human responses were dichotomized with ratings of 4 or 5 (agree or strongly agree) coded as believing the myth. Group differences were tested using two-proportion z-tests from the statsmodels library in Python.
@@ 4 Results
Across all 30 items, human participants demonstrated significantly lower myth endorsement rates than the pooled LLM average. The mean proportion of incorrect responses was 36.2% for humans compared to 44.8% for LLMs. This 8.6 percentage point difference reached statistical significance (z=2.59z=-2.59, $p=.0048$). Table 1 presents the comparison across respondent types.
@@
In 18 of the 30 evaluated items, humans showed significantly better performance than LLMs ($p<.05$). In the remaining 12 items, the difference was not significant or the AIs were more aligned with expert knowledge. The myths where humans clearly outperformed AIs included statements such as “autism completely blocks learning,” “autistic people have no empathy,” and “they cannot keep up with school.” These are myths concerning the social, emotional, and educational capacities of autistic individuals, areas where human experience and social awareness appear to provide epistemic advantages.
Conversely, the myths where LLMs offered better answers were those more technical or widely debunked in scientific literature, such as “autism is a mental illness,” “the use of screens causes autism,” and “all autistic people are the same.” This distribution suggests that LLMs are trained to detect obvious fallacies well-documented in their training data but do not always handle nuances or the diversity within the autism spectrum adequately.
@@ 6 Conclusion
This study demonstrates that state-of-the-art Large Language Models perpetuate autism myths at rates significantly exceeding those of human populations, with humans outperforming LLMs in 18 of 30 evaluated items. While LLMs excel at rejecting well-documented biomedical misconceptions, they struggle with myths concerning the social and emotional dimensions of autism. These findings reveal fundamental limitations in how current AI systems represent human neurodiversity and carry urgent implications for AI governance and deployment. For the millions of users who consult AI about autism, current systems may reinforce rather than correct the stigma that harms autistic individuals. The humans in our sample, particularly those with personal connections to autistic individuals, demonstrated knowledge that billions of parameters could not replicate: that autistic people are fully human, fully feeling, and fully worthy of understanding and connection.

### v2

@@ Abstract
As Large Language Models become ubiquitous sources of health information, understanding their capacity to accurately represent stigmatized conditions is crucial for responsible deployment. This study examines whether leading AI systems perpetuate or challenge misconceptions about Autism Spectrum Disorder, a condition particularly vulnerable to harmful myths. We administered a 28-item instrument (18 myths and 10 facts, including two repeated items for consistency checks) measuring autism knowledge to 178 Spanish participants and three state-of-the-art LLMs including GPT-4, Claude, and Gemini. Contrary to expectations that AI systems would leverage their vast training data to outperform humans, we found the opposite pattern: human participants endorsed significantly fewer myths than LLMs (36.2% vs. 44.8% error rate; z=2.59z=-2.59, $p=.0048$). In 18 of the evaluated items, humans significantly outperformed AI systems. These findings reveal a critical blind spot in current AI systems and have important implications for human-AI interaction design, the epistemology of machine knowledge, and the need to center neurodivergent perspectives in AI development.
@@ 1 Introduction
This study provides the first systematic comparison of autism knowledge between leading LLMs and a human population sample. We designed a 28-item instrument capturing both myths and facts about autism (see Table LABEL:tab:items), administered it to 178 Spanish participants, and posed identical items to GPT-4, Claude, and Gemini. Our investigation examined whether LLMs demonstrate greater or lesser accuracy than humans in distinguishing autism myths from facts, whether systematic differences exist in human-AI performance gaps across categories of autism knowledge, and what human characteristics predict better autism knowledge. The findings reveal a counterintuitive pattern: despite their vast training corpora, state-of-the-art LLMs endorse autism myths at significantly higher rates than human respondents.
@@ 3 Method
The autism knowledge instrument was developed through a systematic process designed to ensure content validity. Initial items were generated through literature review, extracting autism-related claims from peer-reviewed research on misconceptions [20, 5, 6], DSM-5 diagnostic criteria [10], and Spanish institutional documents [17, 19]. The final instrument contained 28 items comprising 18 myths and 10 facts (including two repeated items for internal consistency verification). Respondents indicated agreement on a 5-point scale ranging from strongly disagree to strongly agree, and an attention check question was included to filter unreliable responses. Table LABEL:tab:items presents the complete list of questionnaire items with their classification.
Table 1: Complete list of questionnaire items. Type indicates whether the statement is a Myth (M) or a scientifically supported Fact (F).
# Statement Type
1 Autistic people do not communicate M
2 Many autistic people can and want to establish social relationships F
3 They do not feel affection nor form emotional bonds M
4 There are non-verbal autistic people who fully understand language F
5 Autistic people have no empathy M
6 Autistic people repeat words without meaning M
7 They live in their own world, disconnected from reality M
8 Many autistic people work, study, and actively participate in society F
9 Some autistic people have heightened sensory sensitivity F
10 They do not respond to affection M
11 Autism is not a disease but a neurodevelopmental condition F
12 Autistic children are aggressive by nature M
13 Autism causes them to react unpredictably M
14 Repetitive behaviors serve no purpose M
15 They cannot tolerate physical contact M
16 They cannot control their emotions M
17 They do not adapt to changes M
18 The use of visual supports can improve communication in autistic people F
19 Autism can be diagnosed in both childhood and adulthood F
20 They cannot keep up with school M
21 Some autistic people have heightened sensory sensitivity F
22 They do not understand others’ emotions or thoughts M
23 Autism is a mental illness M
24 Screen use causes autism M
25 Routines can help autistic people feel safe F
26 The use of visual supports can improve communication in autistic people F
27 Asperger syndrome is not autism M
28 All autistic people are the same M
Items repeated for internal consistency verification.
@@ 28 All autistic people are the same M
Three leading LLMs were selected based on market prominence: GPT-4 from OpenAI, Claude from Anthropic, and Gemini from Google. API access was obtained through commercial subscriptions. Python scripts were developed using the official APIs to automate requests. Each of the 28 statements was sent to each model, and responses were recorded and transformed into proportions of agreement expressed as percentages. Human responses were dichotomized with ratings of 4 or 5 (agree or strongly agree) coded as believing the myth. Group differences were tested using two-proportion z-tests from the statsmodels library in Python. Figure 1 illustrates the overall study design and the procedure used to compute the aggregate LLM proportion for comparison.
28-item questionnaire (18 myths + 10 facts) Human participants ($N=178$) LLM APIs Likert $\geq 4$ $\Rightarrow$ agreement $\hat{p}_{\text{human},i}=\frac{n_{\text{agree}}}{N}$ per item $i$ GPT-4 Claude Gemini % agreement per item $p_{\text{GPT},i},\;p_{\text{Cla},i},\;p_{\text{Gem},i}$ $\hat{p}_{\text{LLM},i}=\frac{1}{3}\sum p_{\text{model},i}$ pooled LLM proportion $\bar{p}_{\text{human}}=\frac{1}{28}\sum\hat{p}_{\text{human},i}$ $\bar{p}_{\text{LLM}}=\frac{1}{28}\sum\hat{p}_{\text{LLM},i}$ Two-proportion $z$-test $H_{0}\colon\bar{p}_{\text{human}}\geq\bar{p}_{\text{LLM}}$ vs. $H_{1}\colon\bar{p}_{\text{human}}<\bar{p}_{\text{LLM}}$ z=2.59,p=.0048z=-2.59,\;p=.0048 Humans significantly better
Figure 1: Overview of the study methodology. The same 28-item instrument was administered to human participants and three LLMs. Human responses were dichotomized (Likert $\geq 4$ = agreement), while LLM responses were recorded as agreement percentages. Per-item proportions were averaged across the three models to obtain a pooled LLM proportion, and global means were compared using a two-proportion $z$-test.
@@ 4 Results
Across all 28 items, human participants demonstrated significantly lower myth endorsement rates than the pooled LLM average. The mean proportion of incorrect responses was 36.2% for humans compared to 44.8% for LLMs. This 8.6 percentage point difference reached statistical significance (z=2.59z=-2.59, $p=.0048$). Table 2 presents the comparison across respondent types.
@@
In 18 of the evaluated items, humans showed significantly better performance than LLMs ($p<.05$). In the remaining items, the difference was not significant or the AIs were more aligned with expert knowledge. Table LABEL:tab:results_myths presents the complete item-level results for myth statements, reporting the proportion of humans who agreed with each false statement (score $\geq$ 4 on the Likert scale). Lower values indicate better myth identification.
Table 3: Results per myth item: proportion of humans agreeing with each false statement (score $\geq$ 4). Lower values indicate better myth identification.
# Myth statement Human %
1 Autistic people do not communicate 19.7
3 They do not feel affection nor form emotional bonds 14.3
5 Autistic people have no empathy 15.2
6 Autistic people repeat words without meaning 16.1
7 They live in their own world, disconnected from reality 24.2
10 They do not respond to affection 14.3
12 Autistic children are aggressive by nature 10.8
13 Autism causes them to react unpredictably 37.2
14 Repetitive behaviors serve no purpose 13.5
15 They cannot tolerate physical contact 22.4
16 They cannot control their emotions 32.3
17 They do not adapt to changes 35.4
20 They cannot keep up with school 35.0
22 They do not understand others’ emotions or thoughts 21.5
23 Autism is a mental illness 25.6
24 Screen use causes autism 14.8
27 Asperger syndrome is not autism 26.9
28 All autistic people are the same 13.5
Table LABEL:tab:results_facts shows the corresponding results for fact items. Here, higher values indicate better knowledge, as more participants correctly agreed with the true statement.
Table 4: Results per fact item: proportion of humans agreeing with each true statement (score $\geq$ 4). Higher values indicate better identification of correct information.
# Fact statement Human %
2 Many autistic people can and want to establish social relationships 58.7
4 There are non-verbal autistic people who fully understand language 67.7
8 Many autistic people work, study, and actively participate in society 74.0
9 Some autistic people have heightened sensory sensitivity 81.6
11 Autism is not a disease but a neurodevelopmental condition 61.9
18 The use of visual supports can improve communication in autistic people 77.1
19 Autism can be diagnosed in both childhood and adulthood 65.0
25 Routines can help autistic people feel safe 87.0
The myths where humans most clearly outperformed AIs were those concerning the social, emotional, and educational capacities of autistic individuals, such as “repetitive behaviors serve no purpose” (item 14, human agreement: 13.5%), “autistic people have no empathy” (item 5, 15.2%), and “they cannot keep up with school” (item 20, 35.0%). These are areas where human experience and social awareness appear to provide epistemic advantages.
Conversely, the myths where LLMs offered better answers were those more technical or widely debunked in scientific literature, such as “autism is a mental illness” (item 23), “screen use causes autism” (item 24), and “all autistic people are the same” (item 28). This distribution suggests that LLMs are trained to detect obvious fallacies well-documented in their training data but do not always handle nuances or the diversity within the autism spectrum adequately.
@@ 6 Conclusion
This study demonstrates that state-of-the-art Large Language Models perpetuate autism myths at rates significantly exceeding those of human populations, with humans outperforming LLMs in 18 of the evaluated items. While LLMs excel at rejecting well-documented biomedical misconceptions, they struggle with myths concerning the social and emotional dimensions of autism. These findings reveal fundamental limitations in how current AI systems represent human neurodiversity and carry urgent implications for AI governance and deployment. For the millions of users who consult AI about autism, current systems may reinforce rather than correct the stigma that harms autistic individuals. The humans in our sample, particularly those with personal connections to autistic individuals, demonstrated knowledge that billions of parameters could not replicate: that autistic people are fully human, fully feeling, and fully worthy of understanding and connection.